(a) According to the Ministry of Health, the height of Guinean travellers who were quarantined in Tamale for the novel coronavirus were normally distributed about a mean of 160cm and a standard deviation of 8cm. Find the probability that a traveller selected at random has (i) Height between 148cm and 175cm (ii) Height above 164cm
(iii) Height below 179cm
1
Expert's answer
2020-06-07T14:55:20-0400

Let X = the random variable denoting the height of a Guinean traveller who was quarantined in Tamale for the novel corona virus


Then by the problem,


X ~ N("\mu" = 160, "\sigma"2 = 82)


Therefore,


Z = "\frac{X-\mu}{\sigma}" ~ N(0, 1), Z is the standard normal variate


(i) The probability that a traveler selected at random has height between 148 cm and 175 cm


= P(148 "\leq" X "\leq" 175)


= P("\frac{148-160}{8}\leq\frac{X-160}{8}\leq\frac{175-160}{8}")


= P(-1.5 "\leq" Z "\leq" 1.88)


= "\Phi"(1.88) - "\Phi"(-1.5)


= 0.9699 - 0.0668


= 0.9031


Answer: The probability that a traveler selected at random has height between 148 cm and 175 cm is 0.9031.


(ii) The probability that a traveler selected at random has height above 164 cm


P(X ">" 164)


= P("\frac{X-160}{8}>\frac{164-160}{8}")


= P(Z ">" 0.5)


= 1 - P(Z "\leq" 0.5)


= 1 - "\Phi"(0.5)


= 1 - 0.6915


= 0.3085


Answer: The probability that a traveler selected at random has height above 164 cm 0.3085.


(iii) The probability that a traveler selected at random has height below 179 cm


P(X < 179)


= P("\frac{X-160}{8}<\frac{179-160}{8}")


= P(Z < 2.38)


= "\Phi"(2.38)


= 0.9913


Answer: The probability that a traveler selected at random has height below 179 cm is 0.9913.